# Cost-Efficient Continual Learning with Sufficient Exemplar Memory

### v1

@@ 1 Introduction
  • We propose a novel weight-space manipulation strategy that achives superior performance while drastically reducing computational overhead compared to state-of-the-art algorithms.
@@ 2 Related Works
Continual learning   Continual learning (CL) has been actively studied in various scenarios and methodological categories. Among the three scenarios of CL (Van de Ven & Tolias, 2019), class-incremental learning (class-IL) has been considered the most challenging and has been the most actively studied scenario (Masana et al., 2022). CL algorithms (including class-IL) can be categorized into regularization-based approaches, which penalize changes to important parameters for past tasks (Kirkpatrick et al., 2017; Aljundi et al., 2018; Cha et al., 2020), rehearsal-based approaches, which store and replay exemplars from past tasks (Rebuffi et al., 2001; Cha et al., 2023), and expansion-based approaches, which expand the model’s capacity to balance the trade-off between stability and plasticity (Yan et al., 2021; Wang et al., 2022). Additional approaches focus on addressing classifier bias toward recent tasks while using the exemplars (Wu et al., 2019; Zhao et al., 2020). While exemplar-based methods have demonstrated state-of-the-art performance, they typically rely on strict memory constraints, often limiting memory size to a small percentage of the dataset. Recent studies challenge the necessity of these strict memory constraints, highlighting that the computational cost of maintaining and processing memory—especially GPU usage—can far outweigh storage costs (Prabhu et al., 2023; Chavan et al., 2023). This shift in perspective opens the door to relaxing memory limits in order to reduce training costs, which is the focus of our work.
@@ 5 Experiment
Experimental settings   We evaluate our algorithm on two standard class-incremental learning (class-IL) benchmarks (Masana et al., 2022) using the PyCIL framework (Zhou et al., 2023). CIFAR100 is an image classification dataset with 100 classes, which is split into 10 sequential tasks, each containing 10 classes. ImageNet100 is another classification dataset with 100 classes, which is also split into 10 sequential tasks, each containing 10 classes. We compare our approach with six exemplar-based baselines, including iCaRL (Rebuffi et al., 2001), BiC (Wu et al., 2019), WA (Zhao et al., 2020), DER (Yan et al., 2021), and FOSTER (Wang et al., 2022). Additionally, we include Replay, a naive baseline that finetunes the model using both the current dataset and stored exemplars. All experiments use ResNet-32 for CIFAR-100 and ResNet-18 for ImageNet-100 as the backbone model (He et al., 2016). We conduct experiments across five random seeds and report the final class-IL accuracy, computed as the average classification accuracy across all tasks on the validation datasets after training on the final task (Masana et al., 2022). Lastly, the experimental details are reported in Section A.5
@@ A.4 Experimental Results and Analysis
Memory Size Method 20 40 60 80 100 120 140 160 180 200 220 240 260 280 300 350 400 450 500 DER 64.30 67.56 69.64 70.66 71.68 73.05 73.37 73.41 73.52 74.02 74.85 75.15 74.74 74.96 74.80 75.56 75.44 75.77 75.55 FOSTER 65.83 64.95 65.36 67.00 68.67 69.85 70.84 72.48 73.33 73.46 73.87 74.63 75.16 75.96 76.01 76.88 77.39 77.67 78.24 Replay 48.92 56.59 60.78 63.30 65.12 66.67 68.00 69.23 70.10 70.84 71.61 72.16 73.10 73.19 73.46 74.643 75.38 75.85 76.90 iCaRL 49.56 57.74 61.57 64.14 65.91 67.54 69.21 69.86 70.43 71.42 71.71 72.31 72.83 73.57 73.91 74.85 75.28 76.19 76.78 BiC 54.28 60.48 62.47 63.63 65.84 67.18 67.87 68.19 68.77 68.97 69.46 69.72 70.15 70.34 70.85 71.67 72.16 72.95 72.68 WA 60.87 64.15 65.36 66.47 67.82 68.42 69.09 69.85 70.13 70.40 70.57 71.04 71.11 71.24 71.27 71.92 72.61 73.35 73.52 Ours (w/o avg.) 50.23 58.19 62.12 65.01 66.50 68.25 69.69 70.85 71.66 72.33 72.64 73.78 74.1 74.81 75.00 75.94 76.98 77.37 77.92 Ours (w/o reset) 48.73 56.89 60.81 63.43 65.22 66.50 68.28 68.84 69.85 70.81 71.29 72.03 72.77 73.50 73.49 74.68 74.99 75.57 76.47 Ours 51.91 59.66 63.88 66.24 68.41 69.96 70.92 71.99 72.67 73.48 74.11 74.8 74.87 75.65 75.83 76.89 77.44 77.95 78.25

### v2

@@ 1 Introduction
  • We propose a novel weight-space manipulation strategy that achieves superior performance while drastically reducing computational overhead compared to state-of-the-art algorithms.
@@ 2 Related Works
Continual learning   Continual learning (CL) has been actively studied in various scenarios and methodological categories. Among the three scenarios of CL (Van de Ven & Tolias, 2019), class-incremental learning (class-IL) has been considered the most challenging and has been the most actively studied scenario (Masana et al., 2022). CL algorithms (including class-IL) can be categorized into regularization-based approaches, which penalize changes to important parameters for past tasks (Kirkpatrick et al., 2017; Aljundi et al., 2018; Cha et al., 2020), rehearsal-based approaches, which store and replay exemplars from past tasks (Rebuffi et al., 2001; Cha et al., 2023), and expansion-based approaches, which expand the model’s capacity to balance the trade-off between stability and plasticity (Yan et al., 2021; Wang et al., 2022). Additional approaches focus on addressing classifier bias toward recent tasks while using the exemplars (Wu et al., 2019; Zhao et al., 2020). While exemplar-based methods have demonstrated state-of-the-art performance, they typically rely on strict memory constraints, often limiting memory size to a small percentage of the dataset. Recent studies challenge the necessity of these strict memory constraints, highlighting that the computational cost of maintaining and processing memory—especially GPU usage—can far outweigh storage costs (Prabhu et al., 2023; Chavan et al., 2023; Harun et al., 2023). This shift in perspective opens the door to relaxing memory limits in order to reduce training costs, which is the focus of our work.
@@ 5 Experiment
Experimental settings   We evaluate our algorithm on two standard class-incremental learning (class-IL) benchmarks (Masana et al., 2022) using the PyCIL framework (Zhou et al., 2023). CIFAR100 is an image classification dataset with 100 classes, which is split into 10 sequential tasks, each containing 10 classes. ImageNet100 is another classification dataset with 100 classes, which is also split into 10 sequential tasks, each containing 10 classes. We compare our approach with six exemplar-based baselines, including iCaRL (Rebuffi et al., 2001), BiC (Wu et al., 2019), WA (Zhao et al., 2020), DER (Yan et al., 2021), and FOSTER (Wang et al., 2022). Additionally, we include Replay, a naive baseline that finetunes the model using both the current dataset and stored exemplars. All experiments use ResNet-32 for CIFAR-100 and ResNet-18 for ImageNet-100 as the backbone  (He et al., 2016). We conduct experiments across five random seeds and report the final class-IL accuracy, computed as the average classification accuracy across all tasks on the validation datasets after training on the final task (Masana et al., 2022). Experimental details are reported in Section A.5
@@ A.4 Experimental Results and Analysis
Memory Size Method 20 40 60 80 100 120 140 160 180 200 220 240 260 280 300 350 400 450 500 DER 64.30 67.56 69.64 70.66 71.68 73.05 73.37 73.41 73.52 74.02 74.85 75.15 74.74 74.96 74.80 75.56 75.44 75.77 75.55 FOSTER 65.83 64.95 65.36 67.00 68.67 69.85 70.84 72.48 73.33 73.46 73.87 74.63 75.16 75.96 76.01 76.88 77.39 77.67 78.24 Replay 48.92 56.59 60.78 63.30 65.12 66.67 68.00 69.23 70.10 70.84 71.61 72.16 73.10 73.19 73.46 74.643 75.38 75.85 76.90 iCaRL 49.56 57.74 61.57 64.14 65.91 67.54 69.21 69.86 70.43 71.42 71.71 72.31 72.83 73.57 73.91 74.85 75.28 76.19 76.78 BiC 54.28 60.48 62.47 63.63 65.84 67.18 67.87 68.19 68.77 68.97 69.46 69.72 70.15 70.34 70.85 71.67 72.16 72.95 72.68 WA 60.87 64.15 65.36 66.47 67.82 68.42 69.09 69.85 70.13 70.40 70.57 71.04 71.11 71.24 71.27 71.92 72.61 73.35 73.52 Ours (w/o avg.) 50.23 58.19 62.12 65.01 66.50 68.25 69.69 70.85 71.66 72.33 72.64 73.78 74.1 74.81 75.00 75.94 76.98 77.37 77.92 Ours (w/o reset) 48.73 56.89 60.81 63.43 65.22 66.50 68.28 68.84 69.85 70.81 71.29 72.03 72.77 73.50 73.49 74.68 74.99 75.57 76.47 Ours 51.91 59.66 63.88 66.24 68.41 69.96 70.92 71.99 72.67 73.48 74.11 74.80 74.87 75.65 75.83 76.89 77.44 77.95 78.25